OpenStudy (anonymous):

For i = , if 3i (2 + 5i) = x + 6i, then x = ? Explain How.

OpenStudy (smurfy14):

x=-15

OpenStudy (anonymous):

3i(2+5i) = 6i + 15i^2 = 6i + 15*(-1) = 6i - 15. So if this is x + 6i, the x must be the -15.

OpenStudy (smurfy14):

i^2=-1 so when you distribute and get 15i^2 it reallt is -15 and 6i cancels so x=-15

OpenStudy (anonymous):

yes